OpenStudy (anonymous):

Integrate (e^(x^2)) / (x^3). By looking at this I'm assuming there is going to be a u substitution followed by an integration by parts. I'm having trouble picking a good u value that will make the integration by parts simpler.

OpenStudy (amistre64):

i dont thnk e^(x^2) can be integrated so that would have to account for the "u" part i believe

OpenStudy (amistre64):

at some point I believe youll get an integral on the right that matches your original to work with

OpenStudy (amistre64):

\[\int \frac{e^{x^2}}{x^3}dx=-\frac{e^{x^2}}{2x^2}-\int \frac{2x\ e^{x^2}}{-2x^2} dx\] \[\int \frac{e^{x^2}}{x^3}dx=-\frac{e^{x^2}}{2x^2}+\int \frac{\ e^{x^2}}{x} dx\] hmmmm

OpenStudy (amistre64):

the wolfram give a Ei function but i got no idea what that represents

OpenStudy (anonymous):

The Ei function on wolfram alpha has something to do with e^(x^2) not being able to be integrated I believe. It makes sense to make that "u". Then dv would have to be (x^-3). V = the integral of (x^-3) which is what you showed I believe. Thank you for your help.

OpenStudy (amistre64):

youre welcome, good luck :)